The displacement associated with a wave is given by y(x,t)=0.1 cos(0.2x-2t)
Where x and y are measured in centimeters and t in seconds, the frequency and the velocity are
respectively

Grade:12th pass

1 Answers

ats
3 years ago
The displacement of the wave is given by the equation given in the question.
From this equation we can find that, 
                       k = 0.2 cm-1 and w = 2
Wave velocity = coeffecient of t / coeffecient of x 
                       = 2 / 0.2
                       = 10 cm/s
Frequency = w / 2 pi
                  = 1/ pi s-1     where pi = 3.14
 
Hope it helps! Pls approve this answer if you found it useful.
All the best!